# London WC2R 2LS, UK

### v1

@@ 4.3.1 KT and CYT geometries and conformal invariance
is the curvature of $\lambda$. Using these definitions, the last line of (110) can be verified by a direct calculation. A consequence of (110) is
@@
Some further simplification is possible upon a more careful choice of the frame $(\lambda^{\alpha},\mathbf{e}^{i};\alpha=1,\dots,2p,i=1,\dots 2n-2p)$. For this notice that $\iota_{\widehat{V}_{\alpha}}G^{\perp}=0$ and ${\mathcal{L}}_{\widehat{V}_{\alpha}}G^{\perp}=0$. Thus, $G^{\perp}$ is orthogonal to the directions spanned by the vector fields $\widehat{V}_{\alpha}$ and $\widehat{V}^{K}\partial_{K}G^{\perp}_{IJ}=0$. As a result, one can always choose a frame $\mathbf{e}^{i}$ such that
$V^{J}_{a}\partial_{J}\mathbf{e}_{I}^{i}=0~{}.$ (116)
@@
$\widehat{R}_{\alpha\beta km}=H_{\alpha\beta\gamma}{\mathcal{F}}^{\gamma}_{km}-{\mathcal{F}}_{\alpha ik}{\mathcal{F}}_{\beta j}{}^{k}+{\mathcal{F}}_{\beta ik}{\mathcal{F}}_{\alpha j}{}^{k}~{},$ (119)
@@
^$\widehat{\nabla}_{j}^{\perp}{\mathcal{F}}^{\beta}_{km}=\partial_{j}{\mathcal{F}}^{\beta}_{km}-\widehat{\Omega}_{j}{}^{n}{}_{k}{\mathcal{F}}^{\beta}_{nm}--\widehat{\Omega}_{j}{}^{n}{}_{m}{\mathcal{F}}^{\beta}_{kn}~{},$ (120)
@@
The decomposition of the geometry in (110) and the subsequent analysis presented applies to all manifolds $M^{D}$ that admit a connection with skew-symmetric torsion $\widehat{\nabla}$ and $\widehat{\nabla}$-covariantly constant vector fields. Turning to KT manifolds $M^{2n}$ a consequence of (107) is that the Hermitian form $I$ can be decomposed as
@@
where $I_{\alpha\beta}$ are constants because of ^$\widehat{\nabla}I=0$ and $I^{\perp}$ is again orthogonal to all directions spanned by the vector fields generated by $\mathfrak{g}$ on $M^{2n}$. As it has already been mentioned the integrability of the complex structure $I$ implies that $H$ is a $(2,1)\oplus(1,2)$ form on $M^{D}$. In terms of this frame this condition decomposes as
@@
Furthermore, the Lie derivative of the Hermitian form along the vector fields $\widehat{V}_{\alpha}$ in the same frame is
@@
It is clear from this that the Hermitian form $I$ is invariant under the action of $\mathfrak{g}$ iff the Hermitian form of the fibre is invariant under the adjoint action $\mathfrak{g}$ and the curvature of $\lambda$ is a (1,1) form. It is well known that the former condition can only be satisfied iff $\mathfrak{g}$ is abelian [^15] One way to see this is to notice that there always exist a compact Lie group $\mathcal{G}$ with Lie algebra $\mathfrak{g}$ as $\mathfrak{g}$ admits a positive definite invariant inner product under the $\mathfrak{g}$-adjoint action. Then, the vanishing of the first condition in (126) implies that the associated Hermitian form on $\mathcal{G}$ is bi-invariant and so covariantly constant with respect to the Levi-Civita connection. Thus $\mathcal{G}$ is Kähler and the only such (compact) group manifolds are tori of even dimension.. Furthermore, the vanishing of the second condition in (126), i.e. the restriction for $\mathcal{F}$ to be a (1,1) form, implies the middle condition for the integrability of the complex structure $I$ in (124). Though, the middle condition for the integrability of the complex structure is weaker than the second condition in (126). This is significant as this allows the construction of examples of KT manifolds for which their Hermitian form is not invariant under a group action – the invariance of the Hermitian form under the action of $\mathfrak{g}$ may not be one of the requirements of the construction. It remains to give the Lee form of the KT geometry. This is given as
@@
To begin let us construct examples of KT manifolds with $\widehat{\nabla}$-covariantly constant vector fields. For this consider principal bundles with fibre a group manifold $\mathcal{G}$ of dimension $2p$ that admits a left invariant KT structure, a principal bundle connection $\lambda$ and base space $N^{2n-2p}$ a KT manifold. In particular, the left invariant KT structure on $\mathcal{G}$ is described by the metric $G^{\mathcal{G}}$, torsion 3-form $H^{\mathcal{G}}$ and Hermitian form $I^{\mathcal{G}}$ as
@@
respectively, where $G^{\mathcal{G}}$ and $H^{\mathcal{G}}$ are bi-invariant, i.e. invariant under both the left and right actions of $\mathcal{G}$ on itself, and $(L^{\alpha},\alpha=1,\dots,2p)$ is a left invariant (co-)frame on $\mathcal{G}$. The torsion $H^{\mathcal{G}}$ is given by the structure constants of $\mathcal{G}$. The components $G^{\perp},H^{\perp}$ and $I^{\perp}$ of the metric, torsion and Hermitian form on $M^{2n}$ are pull-back of those on $N^{2n-2p}$. The Hermitian structure $I$ on $M^{2n}$ can be written as in (121) using the components $I_{\alpha\beta}$ of the Hermitian form on the fibre but now taken with respect to the connection $\lambda$ as well as $I^{\perp}$ from the base space. It turns out that with these data $M^{2n}$ is a KT manifold, i.e. $H=-\iota_{I}{\mathrm{d}}I$, iff the curvature of $\lambda$ is a (1,1)-form on $N^{2n-2p}$. Incidentally, this is also the condition, together with the integrability of the complex structure $I^{\perp}$ on $N^{2n-2p}$ and that on the fibre $\mathcal{G}$, for the complex structure $I$ to be integrable on $M^{2n}$. However, the Hermitian form $I$ is not invariant under the action of $\mathcal{G}$, unless $\mathcal{G}$ is abelian, as the vanishing of first condition in (126) required for invariance cannot be satisfied.
@@
By convention, on principal bundles $\mathcal{G}$ acts on each fibre from the right – from now on we shall denote this action with $\mathcal{G}_{R}$. The metric $G$ and torsion $H$ on $M^{2n}$, as described in the previous paragraph, are invariant under the action of $\mathcal{G}_{R}$. However, the complex structure $I$ may not be invariant under the same action unless the Hermitian form of the fibre $\mathcal{G}$ is bi-invariant. As it has been already been explained this can only happen iff $\mathcal{G}$ is abelian. Another possibility arises whether $M^{2n}$ is a product $M^{2n}=\mathcal{G}\times N^{2n-2p}$ and $\lambda$ is the trivial connection – in particular, ${\mathcal{F}}=0$. In such a case, $\lambda$ can be identified with the left invariant 1-forms on $\mathcal{G}$ and the metric, torsion and complex structure can be constructed as in (110) and (121), respectively. Therefore, the KT structure on $M^{2n}$ is the trivial sum of that of the fibre with that of the base space. Of course $G$ and $H$ are invariant under $\mathcal{G}_{R}$ but $I$ is not. However, in this case $I$ is invariant under the left action $\mathcal{G}_{L}$ on $\mathcal{G}$$G$ and $H$ are also invariant under $\mathcal{G}_{L}$.
@@
The holonomy of $\widehat{\nabla}$ of the principal bundle KT geometries constructed above will be $U(n-p)$. For applications to scale and conformal invariant sigma models, the holonomy of $\widehat{\nabla}$ has to be restricted to $SU(n-p)$. One way to achieve this is to take the fibre of the principal bundle to be a KT group manifold, as in the KT $U(n-p)$ case, and further restrict the curvature of the principal bundle as
@@
i.e. $\lambda$ has to satisfy the $SU(n-p)$ instanton-like condition or equivalently the Hermitian-Einstein condition, see (57), on $N^{2n-2p}$. Moreover, one has to choose the base space $N^{2n-2p}$ to be a CYT manifold. As it can be seen from (119) for such a manifold, the holonomy of $\widehat{\nabla}$ will be included in $SU(n-p)$. Thus $M^{2n}$ will be a CYT manifold. In particular sigma models with such a target space will be scale invariant up to two loops in sigma model perturbation theory.
For sigma models with target space the CYT manifold $M^{2n}$, whose geometry has been described above, to be conformally invariant, the additional condition ^$\widehat{\nabla}(\widehat{\theta}+2{\mathrm{d}}\Phi)=0$ has to be satisfied. The Lee form of $M^{2n}$ can be written as
@@
where $\widehat{\theta}^{\mathcal{G}}$ is the Lee form of $\mathcal{G}$ taken along the vertical directions and $\widehat{\theta}^{\perp}$ is the Lee form of the base space $N^{2n-2p}$. Therefore, the conformal invariance condition ^$\widehat{\nabla}(\widehat{\theta}+2{\mathrm{d}}\Phi)=0$ implies that $N^{2n-2p}$ should satisfy
^$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2{\mathrm{d}}\Phi)=0~{},$ (133)
as $\widehat{\theta}^{\mathcal{G}}$ is $\widehat{\nabla}$-covariantly constant. This is a conformal invariance condition for sigma models with target space $N^{2n-2p}$. For example $N^{2n-2p}$ can be chosen to be conformally balanced. In such a case, the second condition for conformal invariance in (93) will also be satisfied.
More examples can be constructed by again starting with fibre a group manifold with a KT structure and by choosing a connection $\lambda$ to satisfy the instanton-like condition
@@
where $\Lambda$ is a constant [^16] Note that the Bianchi identity (54) together with the restriction of the holonomy of $\widehat{\nabla}$ to be included in $SU(n-p)$ and $\iota_{\widehat{V}_{\alpha}}dH=0$ imply that ${\mathcal{F}}^{\alpha}_{ij}(I^{\perp})^{ij}$ is constant. . This is the familiar Hermitian-Einstein condition on ${\mathcal{F}}$ but now with a “cosmological constant”. Furthermore, the base space $N^{2n-2p}$ can be chosen to be a KT manifold with metric $G^{\perp}$, 3-form torsion $H^{\perp}$ and Hermitian structure $I^{\perp}$. The KT structure on $M^{2n}$ can be constructed from that on $\mathcal{G}$ and $N^{2n-2p}$ using the connection $\lambda$ in the way described in (110) and (121). For $M^{2n}$ to admit a CYT, the Ricci form $\widehat{\rho}$ of the $\widehat{\nabla}$ connection vanish. An inspection of (134) reveals that the Ricci form $\widehat{\rho}^{\perp}$ of the connection $\widehat{\nabla}^{\perp}$ of $N^{2n-2p}$ must satisfy
@@
The first condition implies that the first Chern class of $N^{2n-2p}$ becomes trivial upon pulling it back on $M^{2n}$ while the second condition implies that $\Lambda$ must be in the centre of the Lie algebra of $\mathcal{G}$. The construction of KT and CYT manifolds for $\mathcal{G}$ a torus and base space a Kähler manifold has been proposed before in poon .
@@
where $\widehat{\theta}^{\mathcal{G}}$ and $\widehat{\theta}^{\perp}$ have been defined below (132). Therefore, the conformal invariance condition on $M^{2n}$ reduces to (132) as the remaining components of the Lee form of $M^{2n}$ are $\widehat{\nabla}$-covariantly constant. This allows for more examples as the base space can be chosen to be a conformally balanced KT manifold instead of a conformally balanced CYT manifold, previously required, that is more restrictive.
The above constructions solve all the conditions required for $M^{2n}$ to admit an either KT or CYT structure. However to find examples of CYT manifolds that satisfy the conformal condition, a solution of the condition (12) will also be required [^17] One can find connections for the sigma model gauge sector on $M^{2n}$ whose curvature $F$ satisfies all the necessary conditions. For example, one can pull-back bundles $E$ on $M^{2n}$ from $N^{2n-2p}$ that admit a $SU(n-p)$ instanton-like connection. which can be rewritten using (113) as
@@
This equation can be seen as a condition on the 3-form $H^{\perp}$ on $N^{2n-2p}$ given $\lambda$, $\widetilde{R}$ and $F$.

### v2

@@ 4.3 Geometry of conformally invariant backgrounds with $\widehat{V}\not=0$
is the curvature of $\lambda$. Using these definitions, the last line of (110) can be verified by a direct calculation. A consequence of (110) is that
@@
Some further simplification is possible upon a more careful choice of the frame $(\lambda^{\alpha},\mathbf{e}^{i};\alpha=1,\dots,2p,i=1,\dots 2n-2p)$. For this, notice that $\iota_{\widehat{V}_{\alpha}}G^{\perp}=0$ and ${\mathcal{L}}_{\widehat{V}_{\alpha}}G^{\perp}=0$. Thus, $G^{\perp}$ is orthogonal to the directions spanned by the vector fields $\widehat{V}_{\alpha}$ and $\mathcal{L}_{V_{\alpha}}G^{\perp}=0$. As a result, one can always choose a frame $\mathbf{e}^{i}$ such that
$\mathcal{L}_{V_{\alpha}}\mathbf{e}_{I}^{i}=0~.$ (116)
@@
$\widehat{R}_{\alpha\beta ij}=H_{\alpha\beta\gamma}{\mathcal{F}}^{\gamma}_{ij}-{\mathcal{F}}_{\alpha ik}{\mathcal{F}}_{\beta j}{}^{k}+{\mathcal{F}}_{\beta ik}{\mathcal{F}}_{\alpha j}{}^{k}$~, (119)
@@
$\widehat{\nabla}_{j}^{\perp}{\mathcal{F}}^{\beta}_{km}=\partial_{j}{\mathcal{F}}^{\beta}_{km}-\widehat{\Omega}_{j}{}^{n}{}_{k}{\mathcal{F}}^{\beta}_{nm}--\widehat{\Omega}_{j}{}^{n}{}_{m}{\mathcal{F}}^{\beta}_{kn}~,$ (120)
@@
The decomposition of the geometry in (110) and the subsequent analysis presented applies to all manifolds $M^{D}$ that admit a connection with skew-symmetric torsion $\widehat{\nabla}$ and $^\widehat{\nabla}$-covariantly constant vector fields. Turning to KT manifolds $M^{2n}$ a consequence of (107) is that the Hermitian form $I$ can be decomposed as
@@
where $I_{\alpha\beta}$ are constants because of $\widehat{\nabla}I=0$ and $I^{\perp}$ is again orthogonal to all directions spanned by the vector fields generated by $\mathfrak{g}$ on $M^{2n}$. As it has already been mentioned the integrability of the complex structure $I$ implies that $H$ is a $(2,1)\oplus(1,2)$ form on $M^{D}$. In terms of this frame this condition decomposes as
@@
Furthermore, the Lie derivative of the Hermitian form $I$ along the vector fields $\widehat{V}_{\alpha}$ in the same frame is
@@
It is clear from this that the Hermitian form $I$ is invariant under the action of $\mathfrak{g}$ iff the Hermitian form of the fibre is invariant under the adjoint action $\mathfrak{g}$ and the curvature of $\lambda$ is a (1,1) form. It is well known that the former condition can only be satisfied iff $\mathfrak{g}$ is abelian [^16] One way to see this is to notice that there always exist a compact Lie group $\mathcal{G}$ with Lie algebra $\mathfrak{g}$ as $\mathfrak{g}$ admits a positive definite invariant inner product under the $\mathfrak{g}$-adjoint action. Then, the vanishing of the first condition in (126) implies that the associated Hermitian form on $\mathcal{G}$ is bi-invariant and so covariantly constant with respect to the Levi-Civita connection. Thus $\mathcal{G}$ is Kähler and the only such (compact) group manifolds are tori of even dimension.. Furthermore, the vanishing of the second condition in (126), i.e. the restriction for $\mathcal{F}$ to be a (1,1) form, implies the middle condition for the integrability of the complex structure $I$ in (124). Though, the middle condition for the integrability of the complex structure is weaker than the second condition in (126). This is significant as this allows the construction of examples of KT manifolds for which their Hermitian form is not invariant under a group action [^17] As the vector fields $\widehat{V}_{\alpha}$ we consider are Killing, the Hermitian form $I$ will not be invariant, iff the vector fields $\widehat{V}_{\alpha}$ are not holomorphic. – the invariance of the Hermitian form under the action of $\mathfrak{g}$ may not be one of the requirements of the construction. It remains to give the Lee form of the KT geometry. This is given as
@@
4.3.2 KT and CYT geometries and conformal invariance
Example 1: To begin let us construct examples of KT manifolds with $^\widehat{\nabla}$-covariantly constant vector fields. For this consider principal bundles $M$ with fibre a group manifold $\mathcal{G}$ of dimension $2p$ that admits a left invariant KT structure, a principal bundle connection $\lambda$ and base space $N^{2n-2p}$ a KT manifold. In particular, the left invariant KT structure on $\mathcal{G}$ is described by the metric $G^{\mathcal{G}}$, torsion 3-form $H^{\mathcal{G}}$ and Hermitian form $I^{\mathcal{G}}$ as
@@ 4.3.2 KT and CYT geometries and conformal invariance
respectively, where $G^{\mathcal{G}}$ and $H^{\mathcal{G}}$ are bi-invariant, i.e. invariant under both the left and right actions of $\mathcal{G}$ on itself, and $(L^{\alpha},\alpha=1,\dots,2p)$ is a left invariant (co-)frame on $\mathcal{G}$. The components of the torsion $H^{\mathcal{G}}$ are the structure constants of $\mathcal{G}$. The components $G^{\perp},H^{\perp}$ and $I^{\perp}$ of the metric, torsion and Hermitian form on $M^{2n}$ are pull-back of those on $N^{2n-2p}$. The Hermitian structure $I$ on $M^{2n}$ can be written as in (121) using the components $I_{\alpha\beta}$ of the Hermitian form on the fibre but now taken with respect to the connection $\lambda$ as well as $I^{\perp}$ from the base space. It turns out that with these data $M^{2n}$ is a KT manifold, i.e. $H=-\iota_{I}{\mathrm{d}}I$, iff the curvature of $\lambda$ is a (1,1)-form on $N^{2n-2p}$. Incidentally, this is also the condition, together with the integrability of the complex structure $I^{\perp}$ on $N^{2n-2p}$ and that on the fibre $\mathcal{G}$, for the complex structure $I$ to be integrable on $M^{2n}$. However, the Hermitian form $I$ may not be invariant under the action of $\mathcal{G}$, unless $\mathcal{G}$ is abelian. For non-abelian $\mathcal{G}$, the vanishing of the first condition in (126) required for invariance of $I$ cannot be satisfied.
@@
Before, we proceed further note that by convention, on principal bundles $\mathcal{G}$ acts on each fibre from the right – from now on we shall denote this action with $\mathcal{G}_{R}$. The metric $G$ and torsion $H$ on $M^{2n}$, as described in the previous paragraphs, are invariant under the action of $\mathcal{G}_{R}$. However, the complex structure $I$ may not be invariant under the same action unless the Hermitian form of the fibre $\mathcal{G}$ is bi-invariant. As it has been already been explained this can only happen iff $\mathcal{G}$ is abelian. The holonomy of $\widehat{\nabla}$ of all the principal bundle KT geometries constructed above is U(np)U(n-p). The construction of KT and CYT manifolds for $\mathcal{G}$ a torus and base space a Kähler manifold has been proposed before in EGSP ; poon .
Example 2: Another possibility arises whether $M^{2n}$ is a product $M^{2n}=\mathcal{G}\times N^{2n-2p}$ and $\lambda$ is the trivial connection – in particular, ${\mathcal{F}}=0$. In such a case, $\lambda$ can be identified with the left invariant 1-forms $(L^{\alpha})$ on $\mathcal{G}$ and the metric, torsion and complex structure can be constructed as in (110) and (121), respectively. Therefore, the KT structure on $M^{2n}$ is the trivial sum of that of the fibre with that of the base space. Of course $G$ and $H$ are invariant under $\mathcal{G}_{R}$ but $I$ is not. However, in this case $I$ is invariant under the left action $\mathcal{G}_{L}$ on $\mathcal{G}$$G$ and $H$ are also invariant under $\mathcal{G}_{L}$.
@@
Example 3: For applications to scale and conformal invariant sigma models, the holonomy of $\widehat{\nabla}$ has to be restricted to SU(np)SU(n-p). One way to achieve this is to take the fibre of the principal bundle to be a KT group manifold, as in the first KT example above, and further restrict the curvature of the principal bundle as
@@
i.e. $\lambda$ has to satisfy the SU(np)SU(n-p) instanton-like condition or equivalently the Hermitian-Einstein condition, see (57), on $N^{2n-2p}$. Moreover, one has to choose the base space $N^{2n-2p}$ to be a CYT manifold. As it can be seen from (119), for such a manifold the holonomy of $\widehat{\nabla}$ will be included in SU(np)SU(n-p). Thus $M^{2n}$ will be a CYT manifold. In particular sigma models with such a target space will be scale invariant up to two loops in sigma model perturbation theory.
For sigma models with target space the CYT manifold $M^{2n}$, whose geometry has been described above, to be conformally invariant, the additional condition $\widehat{\nabla}(\widehat{\theta}+2{\mathrm{d}}\Phi)=0$ has to be satisfied, where the dilaton $\Phi$ is chosen [^18] One can allow for a “linear” dilaton along the fibre as the conformal invariance condition implies ${\mathrm{d}}{\mathcal{L}}_{\widehat{V}_{\alpha}}\Phi=0$. But this additional restriction will suffice for the analysis of the conditions for conformal invariance presented below. such that ${\mathcal{L}}_{\widehat{V}_{\alpha}}\Phi=0$. The Lee form of $M^{2n}$ can be written as
@@
where $\widehat{\theta}^{\mathcal{G}}$ is the Lee form of $\mathcal{G}$ taken along the vertical directions and $\widehat{\theta}^{\perp}$ is the Lee form of the base space $N^{2n-2p}$. Therefore, the conformal invariance condition $\widehat{\nabla}(\widehat{\theta}+2{\mathrm{d}}\Phi)=0$ implies that $N^{2n-2p}$ should satisfy
$\widehat{\nabla}^{\perp}(\widehat{\theta}^{\perp}+2{\mathrm{d}}\Phi)=0~,$ (133)
as $\widehat{\theta}^{\mathcal{G}}$ is $^\widehat{\nabla}$-covariantly constant. This is a conformal invariance condition for sigma models with target space $N^{2n-2p}$. For example $N^{2n-2p}$ can be chosen to be conformally balanced $\widehat{\theta}^{\perp}=-2{\mathrm{d}}\Phi$ or CY with $\widehat{\theta}^{\perp}=0$ and $\Phi$ constant. This will solve the first condition for conformal invariance in (93).
One can also find connections for the sigma model gauge sector on $M^{2n}$ whose curvature $F$ satisfies all the necessary conditions and in particular the second condition for conformal invariance in (93). For example, one can pull-back bundles $E$ on $M^{2n}$ from $N^{2n-2p}$ that admit a SU(np)SU(n-p) instanton-like connection on $N^{2n-2p}$. Such connection will satisfy the required instanton-like condition on $M^{2n}$ and the second condition for conformal invariance in (93).
Example 4: More examples can be constructed by again starting with the KT geometries of example 1 and by choosing a connection $\lambda$ to satisfy the instanton-like condition
@@
where $\Lambda$ is a constant [^19] Note that the Bianchi identity (54) together with the restriction of the holonomy of $\widehat{\nabla}$ to be included in SU(np)SU(n-p) and $\iota_{\widehat{V}_{\alpha}}dH=0$ imply that ${\mathcal{F}}^{\alpha}_{ij}(I^{\perp})^{ij}$ is constant. It turns out that $\Lambda$ has to commute with all other elements of $\mathfrak{g}$. . This is the familiar Hermitian-Einstein condition on $\mathcal{F}$} but now with a “cosmological constant”. Furthermore, the base space $N^{2n-2p}$ can be chosen to be a KT manifold with metric $G^{\perp}$, 3-form torsion $H^{\perp}$ and Hermitian structure $I^{\perp}$. The KT structure on $M^{2n}$ can be constructed from that on $\mathcal{G}$ and $N^{2n-2p}$ using the connection $\lambda$ in the way described in (110) and (121). For $M^{2n}$ to admit a CYT, the Ricci form $\widehat{\rho}$ of the $\widehat{\nabla}$ connection vanish. An inspection of (134) reveals that for this the Ricci form $\widehat{\rho}^{\perp}$ of the connection $\widehat{\nabla}^{\perp}$ of $N^{2n-2p}$ must satisfy
@@
The first condition implies that the first Chern class of $N^{2n-2p}$ becomes trivial upon pulling it back on $M^{2n}$ while the second condition implies that $\Lambda$ must be in the centre of the Lie algebra of $\mathcal{G}$.
@@
where $\widehat{\theta}^{\mathcal{G}}$ and $\widehat{\theta}^{\perp}$ have been defined below (132). Therefore, the conformal invariance condition on $M^{2n}$ reduces to (133) as the remaining components of the Lee form of $M^{2n}$ are $^\widehat{\nabla}$-covariantly constant. This allows for more examples as the base space can be chosen to be for example conformally balanced KT or Kähler manifold instead of a conformally balanced CYT or CY manifold, previously required in example 3, that is more restrictive.
Connections for the sigma model gauge sector that satisfy all the required conditions, including the second condition in (93) required for conformal invariance, can be constructed as described at the end of example 3. The only difference is that the gauge sector bundle $E$ is taken to be over the base manifold $N^{2n-2p}$ that is a KT manifold instead of a CYT one.
The above constructions solve all the conditions required for $M^{2n}$ to admit an either KT or CYT structure. In addition, they solve that conformal invariance conditions as described in (93). However to find examples of CYT manifolds that satisfy all the conformal conditions, i.e. the vanishing conditions of the beta functions, a solution of the condition (12) will also be required for the CYT geometries described in examples 3 and 4. This condition can be rewritten using (113) as
@@
This equation can be seen as a condition [^20] This equation can be solved on $N^{2n-2p}$ in some cases provided that 2n2p=42n-2p=4 and for $N^{4}$ a conformal balanced manifold. For 2n2p>42n-2p>4, this is equation is more challenging and further analysis is needed to search for solutions. on the 3-form $H^{\perp}$ on $N^{2n-2p}$ given $\lambda$, $\widetilde{R}$ and $F$. Because of this, apart from some very special cases, the base space $N^{2n-2p}$ will be required to have non-vanishing torsion $H^{\perp}$.